Exercise 10-07 Sarasota Furniture Company started construction of a combination office and warehouse building for its own use at an estimated cost of $14,500,000 on January 1, 2020. Sarasota expected to complete the building by December 31, 2020. Sarasota has the following debt obligations outstanding during the construction period. Construction loan-12% interest, payable semiannually, issued December 31, 2019 $5,800,000 Short-term loan-10% interest, payable monthly, and principal payable at maturity on May 30, 2021 4,060,000 Long-term loan-11% interest, payable on January 1 of each year. Principal payable on January 1, 2024 2,900,000 Assume that Sarasota completed the office and warehouse building on December 31, 2020, as planned at a total cost of $15,080,000, and the weighted-average amount of accumulated expenditures was $10,440,000. Compute the avoidable interest on this project Avoidable interest : Compute the depreciation expense for the year ended December 31, 2021. Sarasota elected to depreciate the building on a straight-line basis and determined that the asset has a useful life of 30 years and a salvage value of $870,000 Depreciation expense :
Answer #1

Actual Interest

Construction loan = 5,800,000 x 12% = 696,000

Short Term Loan = 4,060,000 x 10% = 406,000

Long-term loan = 2,900,000 x 11% = 319,000

Avoidable Interest

5,800,000 x 12% = 696,000

4,640,000 x 10.417% = 483,333

Total Avoidable interest = 1,179,333

Weighted Average Interest Rate

Short Term Loan = 4,060,000 x 10% = 406,000

Long-term loan = 2,900,000 x 11% = 319,000

Weighted Average Interest Rate = 725,000/6,960,000

=10.417%

Since avoidable interest is lower than actual interest, we will use avoidable interest.

Cost = 15,080,000 + 1,179,333 = 16,259,333

Depreciation = (16,259,333 - 870,000)/30

=512,977.77